# DOPL: Direct Online Preference Learning for Restless Bandits with Preference Feedback

### v1

@@ E.2 Proof of Lemma 7
Since DOPL designs index not directly based on the preference matrix $\mathbf{F}$, but on the preference-reference term $Q_{n}(s)$ as defined in Lemma 2. To characterize the impact on the preference-reference term $Q_{n}(s)$ made by the estimation error of $F$, we first define a general function $q(x):=\ln\frac{x}{1-x}$. We aim to analyze the impact of a small perturbation in $x$, denoted by $\varepsilon$, on $q(x)$ with the aid of Taylor Series Approximation of $q(x)$. This is presented in the following lemma.
@@
$=\sum_{k=1}^{K}\sum_{n=1}^{N}\langle\omega_{n}^{k},Q_{n}(\tilde{\mathbf{F}}^{k})-Q_{n}(\mathbf{F})\rangle$
$=\sum_{k=1}^{K}\sum_{n=1}^{N}\langle\omega_{n}^{k},Q_{n}(\tilde{\mathbf{F}}^{k})-Q_{n}(\hat{\mathbf{F}}^{k})+Q_{n}(\hat{\mathbf{F}}^{k})-Q_{n}(\mathbf{F})\rangle$
$=\sum_{k=1}^{K}\sum_{n=1}^{N}\langle\omega_{n}^{k},Q_{n}(\tilde{\mathbf{F}}^{k})-Q_{n}(\hat{\mathbf{F}}^{k})\rangle+\sum_{k=1}^{K}\sum_{n=1}^{N}\langle\omega_{n}^{k}Q_{n}(\hat{\mathbf{F}}^{k})-Q_{n}(\mathbf{F})\rangle,$ (46)
@@
$Term_{1}$ $\leq\max_{i,j}\frac{1}{\mathbf{F}(i,j)(1-\mathbf{F}(i,j))}\Bigg{(}\sum_{k=1}^{K}\sum_{n=1}^{N}\langle\omega_{n}^{k},\tilde{\mathbf{F}}^{k}(:,(\star-1)|\mathcal{S}|+\sigma_{s_{\star}})-\hat{\mathbf{F}}^{k}(:,(\star-1)|\mathcal{S}|+\sigma_{s_{\star}})\rangle\allowdisplaybreaks$
@@
$Term_{1}$ $\leq\max_{i,j}\frac{2}{\mathbf{F}(i,j)(1-\mathbf{F}(i,j))}\sum_{k=1}^{K}\sum_{n=1}^{N}\langle\omega_{n}^{k},\mathbf{d}^{k}_{n}\rangle,$ (48)
@@
Next, we bound $\sum_{k=1}^{K}\sum_{n=1}^{N}\langle\omega_{n}^{k},\mathbf{d}_{n}^{k}\rangle$ in the following lemma.
@@
$\sum_{k=1}^{K}\sum_{n=1}^{N}\langle\omega_{n}^{k},\mathbf{d}_{n}^{k}\rangle\leq\sqrt{2|\mathcal{S}|}\sqrt{\ln\frac{4|\mathcal{S}||\mathcal{A}|NT}{\epsilon}}\cdot\sqrt{NT}.$
@@
$\sum_{k=1}^{K}\sum_{n=1}^{N}\langle\omega_{n}^{k},\mathbf{d}_{n}^{k}\rangle\leq\sum_{k=1}^{K}\sum_{n=1}^{N}\left\langle\omega_{n}^{k},\left[\sqrt{\frac{1}{\max(2C_{n\star}^{k}(s_{i},s_{\star}),1)}\ln\!\Big{(}\frac{4|\mathcal{S}||\mathcal{A}|N(k\!-\!1)H}{\epsilon}\Big{)}}\right]_{i=1}^{|\mathcal{S}|}\right\rangle.$ (49)
@@
$\sum_{k=1}^{K}\sum_{n=1}^{N}\left\langle\omega_{n}^{k},\left[\sqrt{\frac{1}{\max(2C_{n\star}^{k}(s_{i},s_{\star}),1)}\ln\!\Big{(}\frac{4|\mathcal{S}||\mathcal{A}|N(k\!-\!1)H}{\epsilon}\Big{)}}\right]_{i=1}^{|\mathcal{S}|}\right\rangle\allowdisplaybreaks$
$=\sum_{k=1}^{K}\sum_{n=1}^{N}\sum_{s,a}\left\langle\omega_{n}^{k}(s,a),\sqrt{\frac{1}{\max(2C_{n\star}^{k}(s,s_{\star}),1)}\ln\!\Big{(}\frac{4|\mathcal{S}||\mathcal{A}|N(k\!-\!1)H}{\epsilon}\Big{)}}\right\rangle\allowdisplaybreaks$
$\overset{(a)}{\leq}\sum_{k=1}^{K}\sum_{n=1}^{N}\sum_{s,a}\sqrt{\frac{1}{\max(2C_{n\star}^{k}(s,s_{\star}),1)}\ln\!\Big{(}\frac{4|\mathcal{S}||\mathcal{A}|N(k\!-\!1)H}{\epsilon}\Big{)}}$
$\leq\sum_{k=1}^{K}\sum_{n=1}^{N}\sum_{s,a}\sqrt{\frac{1}{2Z^{k}_{n}(s,a)}\ln\frac{4|\mathcal{S}||\mathcal{A}|NT}{\epsilon}}\allowdisplaybreaks$
${\leq}\sqrt{\ln\frac{4|\mathcal{S}||\mathcal{A}|NT}{\epsilon}}\sum_{n=1}^{N}\sum_{(s,a)}\sqrt{{Z_{n}^{K}(s,a)}}\allowdisplaybreaks$
$\overset{(b)}{\leq}\sqrt{2|\mathcal{S}|}\sqrt{\ln\frac{4|\mathcal{S}||\mathcal{A}|NT}{\epsilon}}\sum_{n=1}^{N}\sqrt{\sum_{(s,a)}{Z_{n}^{K}(s,a)}}\allowdisplaybreaks$
$\overset{(c)}{\leq}\sqrt{2|\mathcal{S}|}\sqrt{\ln\frac{4|\mathcal{S}||\mathcal{A}|NT}{\epsilon}}\cdot\sqrt{NT},$
where (a) follows since $\omega_{n}^{k}$ is a probability measure, (b) follows Cauchy-Schwartz inequality and (c) uses the fact that $\sum_{n=1}^{N}\sum_{(s,a)}Z_{n}^{T}(s,a)\leq NT.$ This completes the proof. ∎
@@ E.3 Proof of Lemma 8
${\leq}\frac{2B}{1-D}\sqrt{\log\frac{4|\mathcal{S}||\mathcal{A}|NT}{\eta}}\sum_{n=1}^{N}\sum_{(s^{\prime},a^{\prime})}\sqrt{{Z_{n}^{T}(s^{\prime},a^{\prime})}}\allowdisplaybreaks$
$\overset{(d)}{\leq}\frac{2\sqrt{2|\mathcal{S}|}B}{1-D}\sqrt{\log\frac{4|\mathcal{S}||\mathcal{A}|NT}{\eta}}\sum_{n=1}^{N}\sqrt{\sum_{(s^{\prime},a^{\prime})}{Z_{n}^{T}(s^{\prime},a^{\prime})}}\allowdisplaybreaks$
$\overset{(e)}{\leq}\frac{2\sqrt{2|\mathcal{S}|}B}{1-D}\sqrt{\log\frac{4|\mathcal{S}||\mathcal{A}|NT}{\eta}}\cdot\sqrt{NT},$
where (a) follows since $\mathbb{E}\left[\sum_{k=1}^{K}\sum_{h=1}^{H}[P_{\tilde{{\Lambda}}^{k}}F_{\tilde{{\Lambda}}^{k}}](\mathbf{s}_{h})-F_{{\tilde{\Lambda}}^{k}}(\mathbf{s}_{h+1})\right]=0,$ (b) follows standard linear algebra manipulation (akbarzadeh2022learning), (c) follows due to the good event occurs, (d) follows Cauchy-Schwartz inequality and (e) uses the fact that $\sum_{n=1}^{N}\sum_{(s,a)}Z_{n}^{T}(s,a)\leq NT.$

### v2

@@ E.2 Proof of Lemma 5
Since DOPL designs index not directly based on the preference matrix $\mathbf{F}$, but on the preference-reference term $Q_{n}(s)$ as defined in Proposition 2. To characterize the impact on the preference-reference term $Q_{n}(s)$ made by the estimation error of $F$, we first define a general function $q(x):=\ln\frac{x}{1-x}$. We aim to analyze the impact of a small perturbation in $x$, denoted by $\varepsilon$, on $q(x)$ with the aid of Taylor Series Approximation of $q(x)$. This is presented in the following lemma.
@@
$=H\cdot\sum_{k=1}^{K}\sum_{n=1}^{N}\langle\omega_{n}^{k},Q_{n}(\tilde{\mathbf{F}}^{k})-Q_{n}(\mathbf{F})\rangle$
$=H\cdot\sum_{k=1}^{K}\sum_{n=1}^{N}\langle\omega_{n}^{k},Q_{n}(\tilde{\mathbf{F}}^{k})-Q_{n}(\hat{\mathbf{F}}^{k})+Q_{n}(\hat{\mathbf{F}}^{k})-Q_{n}(\mathbf{F})\rangle$
$=H\cdot\sum_{k=1}^{K}\sum_{n=1}^{N}\langle\omega_{n}^{k},Q_{n}(\tilde{\mathbf{F}}^{k})-Q_{n}(\hat{\mathbf{F}}^{k})\rangle+H\cdot\sum_{k=1}^{K}\sum_{n=1}^{N}\langle\omega_{n}^{k},Q_{n}(\hat{\mathbf{F}}^{k})-Q_{n}(\mathbf{F})\rangle,$ (46)
@@
$Term_{1}$ $\leq\max_{i,j}\frac{H}{\mathbf{F}(i,j)(1-\mathbf{F}(i,j))}\Bigg{(}\sum_{k=1}^{K}\sum_{n=1}^{N}\langle\omega_{n}^{k},\tilde{\mathbf{F}}^{k}(:,(\star-1)|\mathcal{S}|+\sigma_{s_{\star}})-\hat{\mathbf{F}}^{k}(:,(\star-1)|\mathcal{S}|+\sigma_{s_{\star}})\rangle\allowdisplaybreaks$
@@
$Term_{1}$ $\leq\max_{i,j}\frac{2}{\mathbf{F}(i,j)(1-\mathbf{F}(i,j))}\sum_{k=1}^{K}\sum_{n=1}^{N}\langle H\cdot\omega_{n}^{k},\mathbf{d}^{k}_{n}\rangle,$ (48)
@@
Next, we bound $\sum_{k=1}^{K}\sum_{n=1}^{N}\langle H\cdot\omega_{n}^{k},\mathbf{d}_{n}^{k}\rangle$ in the following lemma.
@@
$\sum_{k=1}^{K}\sum_{n=1}^{N}\langle H\cdot\omega_{n}^{k},\mathbf{d}_{n}^{k}\rangle\leq\sqrt{2|\mathcal{S}|}\sqrt{\ln\frac{4|\mathcal{S}||\mathcal{A}|NT}{\epsilon}}\cdot\sqrt{NT}.$
@@
$\sum_{k=1}^{K}\sum_{n=1}^{N}\langle H\cdot\omega_{n}^{k},\mathbf{d}_{n}^{k}\rangle\leq\sum_{k=1}^{K}\sum_{n=1}^{N}\left\langle H\cdot\omega_{n}^{k},\left[\sqrt{\frac{1}{\max(2C_{n\star}^{k}(s_{i},s_{\star}),1)}\ln\!\Big{(}\frac{4|\mathcal{S}||\mathcal{A}|N(k\!-\!1)H}{\epsilon}\Big{)}}\right]_{i=1}^{|\mathcal{S}|}\right\rangle.$ (49)
@@
$\sum_{k=1}^{K}\sum_{n=1}^{N}\left\langle H\cdot\omega_{n}^{k},\left[\sqrt{\frac{1}{\max(2C_{n\star}^{k}(s_{i},s_{\star}),1)}\ln\!\Big{(}\frac{4|\mathcal{S}||\mathcal{A}|N(k\!-\!1)H}{\epsilon}\Big{)}}\right]_{i=1}^{|\mathcal{S}|}\right\rangle\allowdisplaybreaks$
$=\sum_{k=1}^{K}\sum_{n=1}^{N}\sum_{s,a}\left\langle H\cdot\omega_{n}^{k}(s,a),\sqrt{\frac{1}{\max(2C_{n\star}^{k}(s,s_{\star}),1)}\ln\!\Big{(}\frac{4|\mathcal{S}||\mathcal{A}|N(k\!-\!1)H}{\epsilon}\Big{)}}\right\rangle\allowdisplaybreaks$
$\overset{(a)}{\leq}\sqrt{\ln\frac{4|\mathcal{S}||\mathcal{A}|NT}{\epsilon}}\sum_{t=1}^{T}\sum_{n=1}^{N}\sum_{(s,a)}\mathds{1}(s_{n}(t)=s,a_{n}(t)=a)\sqrt{\frac{1}{2Z_{n}^{t}(s,a)}}\allowdisplaybreaks$
$=\sqrt{\ln\frac{4|\mathcal{S}||\mathcal{A}|NT}{\epsilon}}\sum_{t=1}^{T}\sum_{n=1}^{N}\sum_{(s,a)}\frac{\mathds{1}(s_{n}(t)=s,a_{n}(t)=a)}{\sqrt{{2Z_{n}^{t}(s,a)}}}\allowdisplaybreaks$
$\overset{(b)}{\leq}2\sqrt{\ln\frac{4|\mathcal{S}||\mathcal{A}|NT}{\epsilon}}\sum_{n=1}^{N}\sum_{(s,a)}{\sqrt{{Z_{n}^{T}(s,a)}}}\allowdisplaybreaks$
[…]
$\overset{(d)}{\leq}2\sqrt{\ln\frac{4|\mathcal{S}||\mathcal{A}|NT}{\epsilon}}\sum_{n=1}^{N}{\sqrt{{|\mathcal{S}||\mathcal{A}|}T}}~{}~{}~{}\text{due to}\sum_{s,a}Z_{n}^{T}(s,a)\leq T\allowdisplaybreaks$
$=2N\sqrt{\ln\frac{4|\mathcal{S}||\mathcal{A}|NT}{\epsilon}}{\sqrt{{|\mathcal{S}||\mathcal{A}|}T}}\allowdisplaybreaks$
$\leq 2\sqrt{2}\sqrt{\ln\frac{4|\mathcal{S}||\mathcal{A}|NT}{\epsilon}}{\sqrt{N^{2}{|\mathcal{S}|}T}}~{}~{}~{}\text{due to}|\mathcal{A}|=2\allowdisplaybreaks$ (50)
where (a) follows since $\omega_{n}^{k}$ is a probability measure, $H\cdot\omega_{n}^{k}(s,a)=Z_{n}^{k}(s,a)-Z_{n}^{k-1}(s,a)=\sum_{h=1}^{H}\mathds{1}(s_{n}(k,h)=s,a_{n}(k,h)=a)$, (b) is due to the fact that for any sequence of numbers $w_{1},w_{2},...,w_{T}$ with $0\leq w_{k}$, and define $W_{k}:=\sum_{i=1}^{k}w_{i}$, we have
$\sum_{k=1}^{T}\frac{w_{k}}{\sqrt{W_{k}}}\leq 2\sqrt{2}\sqrt{W_{T}}.$
(c) follows Jensen’s inequality and the fact that $\sum_{s,a}Z_{n}^{T}(s,a)=T$, and (d) uses the fact that $|\mathcal{A}|=2$. This completes the proof. ∎
@@ E.3 Proof of Lemma 6
$\overset{(d)}{\leq}\frac{4\sqrt{2}B}{1-D}\sqrt{\log\frac{4|\mathcal{S}||\mathcal{A}|NT}{\eta}}\cdot\sqrt{N^{2}|\mathcal{S}|T},$
where (a) follows since $\mathbb{E}\left[\sum_{k=1}^{K}\sum_{h=1}^{H}[P_{\tilde{{\Lambda}}^{k}}F_{\tilde{{\Lambda}}^{k}}](\mathbf{s}_{h})-F_{{\tilde{\Lambda}}^{k}}(\mathbf{s}_{h+1})\right]=0,$ (b) follows standard linear algebra manipulation (akbarzadeh2022learning), (c) follows the same reason as (50). ∎